In a lossy dielectric material of relative permittivity 12, the displacement current density is 25 times greater than conduction current at 1Mhz. Find the conductivity of the dielectric?
1
Expert's answer
2020-05-04T12:46:40-0400
"\frac{J}{J_d}=\frac{\sigma}{\omega\epsilon_r\epsilon_0}"

"\frac{1}{25}=\frac{\sigma}{2\pi (10^6)(12)(8.854\cdot10^{-12})}"

"\sigma=2.67\cdot10^{-5}\frac{S}{m}"